### v1

# Simulation of Hypergraph Algorithms with Looped Transformers

Simulation of Hypergraph Algorithms with Looped Transformers
Xiaoyu Li 7.xiaoyu.li@gmail.com. Independent Researcher.    Yingyu Liang yingyul@hku.hk. The University of Hong Kong. yliang@cs.wisc.edu. University of Wisconsin-Madison.    Jiangxuan Long lungchianghsuan@gmail.com. South China University of Technology.    Zhenmei Shi zhmeishi@cs.wisc.edu. University of Wisconsin-Madison.    Zhao Song magic.linuxkde@gmail.com. The Simons Institute for the Theory of Computing at the University of California, Berkeley.    Zhen Zhuang zhuan244@umn.edu. University of Minnesota.
Looped Transformers have shown exceptional capability in simulating traditional graph algorithms, but their application to more complex structures like hypergraphs remains underexplored. Hypergraphs generalize graphs by modeling higher-order relationships among multiple entities, enabling richer representations but introducing significant computational challenges. In this work, we extend the Loop Transformer architecture to simulate hypergraph algorithms efficiently, addressing the gap between neural networks and combinatorial optimization over hypergraphs. In this paper, we extend the Loop Transformer architecture to simulate hypergraph algorithms efficiently, addressing the gap between neural networks and combinatorial optimization over hypergraphs. Specifically, we propose a novel degradation mechanism for reducing hypergraphs to graph representations, enabling the simulation of graph-based algorithms, such as Dijkstra’s shortest path. Furthermore, we introduce a hyperedge-aware encoding scheme to simulate hypergraph-specific algorithms, exemplified by Helly’s algorithm. The paper establishes theoretical guarantees for these simulations, demonstrating the feasibility of processing high-dimensional and combinatorial data using Loop Transformers. This work highlights the potential of Transformers as general-purpose algorithmic solvers for structured data.
1 Introduction
Large Language Models (LLMs) built on the Transformer architecture [61], such as GPT-4o [47], Claude [1], and OpenAI’s latest o1 and o3 [48], demonstrate the remarkable potential to improve our daily lives. They are transforming domains like AI agents [12], search AI [48], AI assistants [46, 74], and conversational AI [31]. As their capabilities expand, LLMs serve as powerful instruments driving the future of technological innovation. A key emergent capability of LLMs is their adeptness at executing traditional algorithms, showcasing exceptional performance in tasks such as sorting large datasets and solving intricate problems. LLMs have shown particular aptitude in processing diverse data structures and types, including graph-structured data [69], tabular formats [67], and temporal sequences [78]. Notably, the Loop Transformer architecture has demonstrated robust performance in simulating various graph algorithms [13], achieving comparable results to traditional graph processing methods. These developments not only advance our theoretical understanding of Transformer-based models’ computational capabilities but also offer practical alternatives for complex algorithmic tasks. This success raises a compelling question:
Can Loop Transformers extend their algorithmic simulation capabilities to the more complex domain like hypergraphs?
Hypergraphs, which extend traditional graphs by permitting edges, called hyperedges, to connect multiple nodes, offer a more natural representation of many real-world datasets and problems. This added complexity enables hypergraphs to capture higher-order relationships between entities, making them particularly useful in a variety of domains where interactions are not limited to pairwise connections. For instance, in combinatorial optimization, hypergraphs are used to model complex systems such as resource allocation [76] and network flow [72], where multiple entities interact simultaneously in a single operation. Similarly, in recommendation systems [68, 66], hypergraphs can represent user-item interactions more flexibly, enabling the modeling of complex, multi-way relationships, such as collaborative filtering or group-based recommendations. In computational biology, hypergraphs are employed to capture intricate biological relationships, such as protein-protein interaction networks or gene expression data [60], where multiple factors simultaneously contribute to the system’s behavior.
Hypergraph algorithms, however, often pose greater computational challenges than graph algorithms due to their richer structure and exponential growth in possible hyperedge combinations. Bridging the gap between hypergraph algorithms and neural network simulation could unlock new paradigms for solving high-dimensional and combinatorially complex problems.
1.1 Our Contribution
In this paper, we extend the results of [13] that demonstrated the ability of Loop Transformers to simulate graph algorithms. We propose and investigate the hypothesis that Loop Transformers, with their iterative processing and representational flexibility, are capable of simulating hypergraph algorithms. By leveraging the inherent capacity of Transformers for parallel processing and multi-head attention, we aim to show that these models can effectively encode and operate over hypergraph structures. In detail, we show that the Loop Transformer can simulate a degradation mechanism (see detail in Section 6) by using 10 layers, where each attention layer includes three attention heads with constant feature dimensions. Also, the Loop Transformer can represent the Helly algorithm (see detail in Section 6) by using 11 attention layers, where each layer includes three attention heads with constant feature dimensions. These findings demonstrate that the Looped Transformer is a general-purpose computational tool capable of bridging neural networks and combinatorial optimization tasks on structured, high-dimensional data.
The contributions of our work are outlined as follows:
  • We design a degradation mechanism (Theorem 6.1) to simulate graph-based algorithms, e.g., Dijkstra, BFS, DFS, on hypergraphs via reduced graph representations within the Loop Transformer framework.
  • We propose a novel encoding scheme to simulate hyperedge-aware algorithms, e.g., Helly’s algorithm (Theorem 6.2), by incorporating hyperedge-specific operations into Transformer’s iterative processes.
Roadmap.
We introduce the works related to our paper in Section 2. In Section 3, we present the basic concepts, e.g., hypergraph and Loop Transformers. In Section 4, we introduce the construction of some basic operations. In Section 5, we present some simulation results based on the basic operation. In Section 6, we detail our main results. In Section 7, we give a conclusion of our paper.
2 Related Work
2.1 Neural Network Execute Algorithms
In recent years, the integration of neural networks with algorithmic execution has gained significant attention, leading to the emergence of a body of work aimed at improving the efficiency and performance of algorithmic tasks through deep learning techniques. This line of research focuses on leveraging the power of neural networks to either approximate traditional algorithms or directly replace them with learned models. [53] established the computational universality of finite recurrent neural networks with sigmoidal activation functions by proving their capability to simulate any Turing Machine. Later, [49] proved that Transformer architectures achieve Turing completeness through hard attention mechanisms, which enable effective computation and access to internal dense representations. Building with the framework of looped transformer, [18] uses looped transformers as the building block to make a programmable computer, showcasing the latent capabilities of Transformer-based neural networks. [13] demonstrated that looped transformers can implement various graph algorithms, including Dijkstra’s shortest path, Breadth-First Search (BFS), Depth-First Search (DFS), and Kosaraju’s algorithm for strongly connected components, through their multi-head attention mechanism. [18, 70] uses Transformers as the building block to build a programmable computer, showcasing the latent capabilities of Transformer-based neural networks. Beyond [18], [41] proves that a looped 23-layer $\mathrm{ReLU-MLP}$ is capable of performing the basic necessary operation of a programmable computer. [19, 9] showed that looped transformers can efficiently do in-context learning by multi-step gradient descent.
2.2 Hypergraphs in Neural Networks
Hypergraphs have recently emerged as a powerful mathematical model for representing complex relationships in data, and they have found promising applications in deep learning. One of the key motivations for using hypergraphs in neural networks is their ability to capture higher-order interactions among entities, which can be more representative of real-world data than traditional pairwise relationships. Several works have explored leveraging hypergraph-based representations to improve the performance of neural architectures in various domains. For instance, [16] introduced a hypergraph-based approach to enhance graph neural networks (GNNs), where the nodes of a graph were represented as hypernodes in a hypergraph. This structure allowed the model to better capture multi-way interactions, significantly improving results on tasks such as node classification and link prediction. In another study, [71] extended convolutional neural networks (CNNs) to work on hypergraphs, demonstrating how hypergraph convolution can be used to aggregate features over hyperedges, yielding more accurate feature representations in image and video processing tasks. Additionally, [6] proposed a novel method for integrating hypergraph attention mechanisms into deep neural networks, allowing the model to focus on the most influential hyperedges for decision-making. This approach was shown to enhance the interpretability of deep learning models and achieve state-of-the-art performance in tasks involving large-scale multi-modal data. In [4], the authors investigated the use of hypergraphs in the context of deep reinforcement learning, where the structure of the hypergraph facilitated more efficient exploration of state-action spaces, leading to faster convergence in complex environments. Recent relevant research includes works on tensor transformers [51, 3, 44, 33, 75], acceleration techniques [64, 63, 56, 36, 50, 59, 22, 57, 52, 62, 29, 23, 24, 35, 27, 7, 30, 37, 34, 26, 42, 40, 28, 39, 2, 11], and other related studies [14, 54, 20, 43, 58, 8, 65, 45, 15, 17, 10, 32, 25, 55, 77, 73, 38].
3 Preliminaries
In Section 3.1, we introduce the fundamental notations used throughout this work. Section 3.2 outlines the concept of simulation, while Section 3.3 provides an overview of essential concepts related to hypergraphs. Lastly, Section 3.4 describes the architecture of the looped transformer.
3.1 Notation
We represent the set $\{1,2,\ldots,n\}$ as $[n]$. For a matrix $A\in\mathbb{R}^{m\times n}$, the $i$-th row is denoted by $A_{i}\in\mathbb{R}^{n}$, and the $j$-th column is represented as $A_{*,j}\in\mathbb{R}^{m}$, where $i\in[m]$ and $j\in[n]$. For $A\in\mathbb{R}^{m\times n}$, the $j$-th entry of the $i$-th row $A_{i}\in\mathbb{R}^{n}$ is denoted by $A_{i,j}\in\mathbb{R}$. The identity matrix of size $d\times d$ is denoted by $I_{d}\in\mathbb{R}^{d\times d}$. The vector ${\bf 0}_{n}$ denotes a length-$n$ vector with all entries equal to zero, while ${\bf 1}_{n}$ denotes a length-$n$ vector with all entries equal to one. The matrix ${\bf 0}_{n\times d}$ represents an $n\times d$ matrix where all entries are zero. The inner product of two vectors $a,b\in\mathbb{R}^{d}$ is expressed as $a^{\top}b$, where $a^{\top}b=\sum_{i=1}^{d}a_{i}b_{i}$.
3.2 Simulation
Definition 3.1 (Simulation, Definition 3.1 on page 3 in [13]).
We define the following:
  • Let $h_{F}:\mathcal{X}\to\mathcal{Y}$ be the function that we want to simulate.
  • Let $h_{T}:\mathcal{X}^{\prime}\to\mathcal{Y}^{\prime}$ be a neural network.
  • Let $g_{e}:\mathcal{X}\to\mathcal{X}^{\prime}$ be an encoding function.
  • Let $g_{d}:\mathcal{Y}^{\prime}\to\mathcal{Y}$ be a decoding function.
We say that a neural network $h_{T}$ can simulate an algorithm step $h_{F}$ if for all input $x\in\mathcal{X}$, $h_{F}(x)=g_{d}(h_{T}(g_{e}(x)))$. We say that a looped transformer can simulate an algorithm if it can simulate each algorithm step.
3.3 Hypergraph
A weighted hypergraph is expressed as $H:=(V,E,w)$, where $w$ is a function assigning a real-valued weight to each hyperedge. The total number of vertices in the hypergraph is $n_{v}:=|V|$, and the total number of hyperedges is $n_{e}:=|E|$. The vertices are assumed to be labeled sequentially from $1$ to $n_{v}$, and the hyperedges are labeled from $1$ to $n_{e}$.
We define the incident matrix of the weighted hypergraph as follows:
Definition 3.2 (Incident matrix of hypergraph).
Let $H=(V,E,w)$ be a weighted hypergraph, where $V=\{v_{1},v_{2},\ldots,v_{n_{v}}\}$ is the set of vertices and $E=\{e_{1},e_{2},\ldots,e_{n_{e}}\}$ is the set of hyperedges, with $n_{v}=|V|$ and $n_{e}=|E|$. Each hyperedge $e_{j}\in E$ is associated with a weight $w(e_{j})\in\mathbb{R}_{+}$ and consists of a subset of vertices from $V$. The incident matrix $A\in\mathbb{R}^{n_{v}\times n_{e}}_{+}$ of $H$ is defined such that the rows correspond to vertices and the columns correspond to hyperedges. For each vertex $v_{i}\in V$ and hyperedge $e_{j}\in E$, the entry $A_{i,j}$ is given by
$A_{i,j}=\begin{cases}w(e_{j})&\mathrm{if}~{}v_{i}\in e_{j},\\ 0&\mathrm{otherwise.}\end{cases}$
In this paper, we use $X\in\mathbb{R}^{K\times d}$ to represent the input matrix, where $d$ is the feature dimension and $K=\max\{n_{v},n_{e}\}+1$ is the row number we need for simulation. To match the dimension of $X$ and incident matrix, we use a padded version of $A$, which is defined as its original entries of $A$ preserved in the bottom-right block:
Definition 3.3 (padded version of incident matrix).
Let incident matrix of hypergraph $A\in\mathbb{R}^{n_{v},n_{e}}_{+}$ be defined in Definition 3.2, let $K\in\mathbb{N}$ safisfies $K\geq\max\{n_{v},n_{e}\}+1$ is the row number of $X$, we define the padded version of incident matrix $A$ as:
  • Part 1. If $n_{e}>n_{v}$, we define:
    $\widetilde{A}:=\begin{bmatrix}0&\mathbf{0}_{n_{e}}^{\top}\\ \mathbf{0}_{n_{v}}&A\\ \mathbf{0}_{K-n_{v}-1}&\mathbf{0}_{(K-n_{v}-1)\times n_{e}}\end{bmatrix}.$
  • Part 2. If $n_{e}<n_{v}$, we define:
    $\widetilde{A}:=\begin{bmatrix}0&\mathbf{0}_{n_{e}}^{\top}&\mathbf{0}_{K-n_{e}-1}^{\top}\\ \mathbf{0}_{n_{v}}&A&\mathbf{0}_{n_{v}\times(K-n_{e}-1)}\end{bmatrix}.$
  • Part 3. If $n_{e}=n_{v}$, we define:
    $\widetilde{A}:=\begin{bmatrix}0&\mathbf{0}_{n_{e}}^{\top}\\ \mathbf{0}_{n_{v}}&A\end{bmatrix}.$
3.4 Looped Transformers
Following the setting of [13], we use standard transformer layer [61] with an additional attention mechanism that incorporates the incident matrix. The additional attention mechanism is defined as follows:
Definition 3.4 (Single-head attention, [13]).
Let $W_{Q},W_{K}\in\mathbb{R}^{d\times d_{a}}$ be the weight matrices of query and key, $W_{V}\in\mathbb{R}^{d\times d}$ be the weight matrix of value, and $\sigma$ be the hardmax [^1] The hardmax, a.k.a., $\arg\max$, is defined by $[\sigma(\Phi)]_{i}:=\sum_{k\in K}e_{k}/|K|$, where $e_{k}$ is the standard basis vector for any $k\in[K]$ and $K=\{k~{}|~{}\Phi_{ik}=\max\{\Phi_{i}\}\}.$ function. We define the single-head attention $\psi^{(i)}$ as
$\psi^{(i)}(X,\widetilde{A}):=\widetilde{A}\sigma(XW^{(i)}_{Q}{W^{(i)}_{K}}^{\top}X^{\top})XW^{(i)}_{V}.$
Remark 3.5.
In this paper, we set $d_{a}=2$.
The $\psi$ function is an essential construction in the definition of multi-head attention:
Definition 3.6 (Multi-head attention, [13]).
Let $\psi$ be defined in Definition 3.4, let $\widetilde{A}$ be defined in Definition 3.3. We define the multi-head attention $\psi^{(i)}$ as
$f_{\mathrm{attn}}(X,\widetilde{A})~{}:=~{}\sum_{i\in M_{A}}\psi^{(i)}(X,\widetilde{A})+\sum_{i\in M_{A^{\top}}}\psi^{(i)}(X,\widetilde{A}^{\top})+\sum_{i\in M}\psi^{(i)}(X,I_{n+1})+X,$
where $M_{A}$, $M_{A^{\top}}$, $M$ are the index set of the attention incorporated the incident matrix, and the attention incorporated the transpose of the incident matrix, and attention heads for the standard attention which is defined in [61].
Remark 3.7.
The total number of attention heads is $|M|+|M_{A}|+|M_{A^{\top}}|$.
For the MLP (Multilayer Perceptron) layer, we have the following definition.
Definition 3.8 (MLP layer, [13]).
Let $\phi$ be the ReLU function, let $W\in\mathbb{R}^{d\times d}$ be the weight of the MLP, where $d$ is the feature dimension of $X$, let $m$ be the number of layers. For $j=[m]$, we define the MLP layer as
$f_{\mathrm{mlp}}(X):=Z^{(m)}W^{(m)}+X,$
where $Z^{(j+1)}:=\phi(Z^{(j)}W^{(j)})$ and $Z^{(1)}:=X$.
Remark 3.9.
In the construction of this paper, we set the number of layers $m=4$.
Remark 3.10.
In the construction of this paper, we set $X\in\mathbb{R}^{K\times d}$, where $K$ is defined in Definition 3.3, $d$ is a constant independent of $K$. The matrix $X$ stores different variables in its different columns. We use $B_{\mathrm{global}}$ to present the column that only the top one scalar is $1$ while the rest are 00. We use $B_{\mathrm{local}}$ to present the column that only the top one scalar is 00 while the rest are $1$. We use $P$ to present the 2 columns of position embedding, where the first column is $\sin(\theta_{i})$, and the second column is $\cos(\theta_{i})$ for $i\in[K-1]$. We use $P_{\mathrm{cur}}$ to present the position embedding of the current vertex or hyperedge.
Combine the definition of multi-head attention and MLP layer, and we show the definition of the transformer layer:
Definition 3.11 (Transformer layer, [13]).
Let $\widetilde{A}$ be defined in Definition 3.3, let $f_{\mathrm{attn}}$ be defined in Definition 3.6, let $f_{\mathrm{mlp}}$ be defined in Definition 3.8. We define the transformer layer as
$f(X,\widetilde{A}):=f_{\mathrm{mlp}}(f_{\mathrm{attn}}(X,\widetilde{A})).$
Definition 3.12 (Multi layer Transformer, [13]).
Let $\widetilde{A}$ be defined in Definition 3.3, let $m=O(1)$ denote the layer of transformers. Let the transformer layer $f$ be defined in 3.11. We define the multi-layer transformer as
$h_{T}(X,\widetilde{A}):=f_{m}\circ f_{m-1}\circ\cdot\circ f_{1}(X,\widetilde{A}).$
Algorithm 1 Looped Transformer, Algorithm 1 in [18]
1:procedure LoopedTransformer($X\in\mathbb{R}^{K\times d},\widetilde{A}\in\mathbb{R}^{K\times K},\mathrm{termination}\in\{0,1\}$)
2:     while $X[0,\mathrm{termination}]=0$ do
3:         $X\leftarrow h_{T}(X,\widetilde{A})$
4:     end while
5:end procedure
Definition 3.13 (Positional Encoding).
Let $\delta$ be the minimum increment angle, and let $\widehat{\delta}$ be its nearest representable approximation. Define the rotation matrix $R_{\widehat{\delta}}\in\mathbb{R}^{2\times 2}$ as
$R_{\widehat{\delta}}=\begin{bmatrix}\cos\widehat{\delta}&-\sin\widehat{\delta}\\ \sin\widehat{\delta}&\cos\widehat{\delta}\end{bmatrix}.$
Initialize the positional encoding with $p_{0}=\begin{pmatrix}0\\ 1\end{pmatrix}\in\mathbb{R}^{2}$. For each node $i\geq 1$, the positional encoding $p_{i}\in\mathbb{R}^{2}$ is defined recursively by
$p_{i}=R_{\widehat{\delta}}^{\top}p_{i-1}.$
The positional encoding for node $i$ is represented as the tuple $(p_{i}^{(1)},p_{i}^{(2)})$, where $p_{i}^{(1)}$ and $p_{i}^{(2)}$ are the first and second components of the vector $p_{i}$, respectively.
Additionally, the maximum number of distinct nodes that can be uniquely encoded is bounded by
$N_{\max}=\left\lfloor\frac{2\pi}{\widehat{\delta}}\right\rfloor,$
which is determined by the precision of $\widehat{\delta}$.
Algorithm 2 Iteration of minimum value
1:$\mathrm{idx}_{\mathrm{cur}},~{}\mathrm{val}_{\mathrm{cur}}\leftarrow 0,~{}0$
2:$\mathrm{idx}_{\mathrm{best}},~{}\mathrm{val}_{\mathrm{best}}\leftarrow 0,~{}\Omega$
3:$\mathrm{visit}_{\min}\leftarrow{\bf 0}_{n_{v}}\|{\bf 1}_{K-1-n_{v}}$
4:$\mathrm{termination}_{\min}\leftarrow 0$triangleright Initialization of minimum value
5:
6:procedure GetMinimumValue($x\in\mathbb{R}^{d}$)
7:     if $\mathrm{termination}_{\min}=1$ then
8:         $\cdots$ $\triangleright$ Re-initialization of minimum value, Lemma 4.1
9:         $\mathrm{termination}_{\min}\leftarrow 0$
10:     end if
11:     $\mathrm{idx}_{\mathrm{cur}}\leftarrow\mathrm{idx}_{\mathrm{cur}}+1$triangleright Increment, Lemma 4.2
12:     $\mathrm{val}_{\mathrm{cur}}\leftarrow x[\mathrm{idx}_{\mathrm{cur}}]$triangleright Read scalar from column, Lemma 4.3
13:     if $\mathrm{val}_{\mathrm{cur}}<\mathrm{val}_{\mathrm{best}}$ $\triangleright$ Compare, Lemma 4.4
14:         $\mathrm{val}_{\mathrm{best}},~{}\mathrm{idx}_{\mathrm{best}}\leftarrow\mathrm{val}_{\mathrm{cur}},~{}\mathrm{idx}_{\mathrm{cur}}$triangleright Update variables, Lemma 4.1
15:     end if
16:     $\mathrm{visit}_{\min}[\mathrm{idx}_{\mathrm{cur}}]\leftarrow 1$triangleright Write scalar to column, Lemma 4.5
17:     $\mathrm{termination}_{\min}\leftarrow\neg(0~{}\mathrm{in}~{}\mathrm{visit}_{\min})$triangleright Trigger termination, Lemma 4.6
18:end procedure
4 Key Operation Implementation
In Section 4.1, the operation of selection is discussed. In Section 4.2, the operation of increment is discussed. In Section 4.3, the operation of reading scalar from column is discussed. In Section 4.4, the operation of comparison is discussed. In Section 4.5, the operation of writing scalar to column is discussed. In Section 4.6, the operation of termination is discussed. In Section 4.7, the operation of reading from incident matrix is discussed. In Section 4.8, the operation of AND is discussed. In Section 4.9, the operation of repeat AND is discussed. In Section 4.10, the operation of repeat addition is discussed.
4.1 Selection
Lemma 4.1 (Selection).
If the following conditions hold:
  • Let the transformer be defined as Definition 3.11.
  • Let $\Omega$ represent the maximum absolute value within a clause.
  • Let $V_{0},V_{1}$ be the index for clause values, where $X[i,V_{0}],X[i,V_{1}]\in[-\Omega,\Omega]$ for $i\in[K]$.
  • Let $C$ be the index for conditions, where $X[i,C]\in\{0,1\}$ for $i\in[K]$.
  • Let $E$ be the index of the target field.
Then we can show that a single-layer transformer can simulate a selection operation, which achieves the following: if $X[i,C]=1$, the value $X[i,V_{1}]$ is written to $X[i,E]$; otherwise, the value $X[i,V_{0}]$ is written to $X[i,E]$, for all $i\in[K]$.
Proof.
Because only the MLP layer is needed, we set all the parameters in the attention layer to 0 while keeping the residual connection. Let $S_{1},~{}S_{2}$ be the index for the scratchpad. We construct the weights in MLP as follows:
$(W^{(1)})_{a,b}$ =$=\begin{cases}1&\mathrm{if}~{}(a,b)\in\{(V_{0},V_{0}),(V_{1},V_{1}),(C,C),(E,E),(B_{\mathrm{global}},B_{\mathrm{global}}),(B_{\mathrm{local}},B_{\mathrm{local}})\};\\ 0&\mathrm{otherwise,}\end{cases}$
$(W^{(2)})_{a,b}$ =$=\begin{cases}1&\mathrm{if}~{}(a,b)\in\{(E,E),(V_{0},S_{1}),(V_{1},S_{2})\};\\ -\Omega&\mathrm{if}~{}(a,b)\in\{(C,S_{1}),(B_{\mathrm{global}},S_{2}),(B_{\mathrm{local}},S_{2})\};\\ \Omega&\mathrm{if}~{}(a,b)=(C,S_{2});\\ 0&\mathrm{otherwise,}\end{cases}$
$(W^{(3)})_{a,b}$ =$=\begin{cases}1&\mathrm{if}~{}(a,b)\in\{(E,E),(S_{1},S_{1}),(S_{2},S_{1})\};\\ 0&\mathrm{otherwise,}\end{cases}~{}(W^{(4)})_{a,b}=\begin{cases}1&\mathrm{if}~{}(a,b)=(E,S_{1});\\ -1&\mathrm{if}~{}(a,b)=(E,E);\\ 0&\mathrm{otherwise.},\end{cases}$
where $W^{(1)}$ is defined as an identity operator, $W^{(2)}$ is defined to perform the selection, $W^{(3)}$ is defined to sum the terms, $W^{(4)}$ is defined to write the term on scarctchpad to column $E$. ∎
4.2 Increment
Lemma 4.2 (Increment).
If the following conditions hold:
  • Let the transformer be defined as Definition 3.11.
  • Let $\widehat{\delta}$ be the nearest representable approximation of the minimum increment angle in position encoding defined in Definition 3.13.
  • Let $C_{1},C_{2}$ be the index for source position embedding, $D$ be the index for target position embedding.
Then we can show that a single-layer transformer can simulate an increment operation, which achieves $X[1,D_{1}]\leftarrow\sin(\arcsin{(X[1,C_{1}])}+\widehat{\delta})$, $X[1,D_{2}]\leftarrow\cos(\arccos{(X[1,C_{2}])}+\widehat{\delta})$.
Proof.
Because only the MLP layer is needed, we set all the parameters in the attention layer to 0 while keeping the residual connection. Let $S_{1},S_{2}$ be the index for the scratchpad. We construct $W^{(1)}$ corresponding to the rotation matrix, which is defined in Definition 3.13. $W^{(2,3)}$ is defined as identity operator, and $W^{(4)}$ is defined to erase the previous value in $X[:,D]$.
$(W^{(1)})_{a,b}$ =$=\begin{cases}\cos(\widehat{\delta})&\mathrm{if}~{}(a,b)\in\{(C_{1},S_{1}),(C_{2},S_{2})\};\\ -\sin(\widehat{\delta})&\mathrm{if}~{}(a,b)=(C_{1},S_{2});\\ \sin(\widehat{\delta})&\mathrm{if}~{}(a,b)=(C_{2},S_{1});\\ 1&\mathrm{if}~{}(a,b)\in\{(D_{1},D_{1}),(D_{2},D_{2})\};\\ 0&\mathrm{otherwise,}\end{cases}$
$(W^{(2,3)})_{a,b}$ =$=\begin{cases}1&\mathrm{if}~{}(a,b)\in\{(D_{1},D_{1}),(D_{2},D_{2}),(S_{1},S_{1}),(S_{2},S_{2})\};\\ 0&\mathrm{otherwise,}\end{cases}$
$(W^{(4)})_{a,b}$ =$=\begin{cases}1&\mathrm{if}~{}(a,b)\in\{(S_{1},D_{1}),(S_{2},D_{2})\};\\ -1&\mathrm{if}~{}(a,b)\in\{(D_{1},D_{1}),(D_{2},D_{2})\};\\ 0&\mathrm{otherwise.}\end{cases}$
4.3 Read Scalar From Column
Lemma 4.3 (Read scalar from column).
If the following conditions hold:
  • Let the transformer be defined as Definition 3.11.
  • Let $\Omega$ represent the maximum absolute value within a clause.
  • Let $C$ be the index for source row, $D$ be the index for source column.
  • Let $E$ be the index for the target column.
Then, we can show that a single-layer transformer can simulate an increment operation, which achieves writing $X[1,E]\leftarrow X[C,D]$.
Proof.
The key purpose is to move the information from the local variable to the global variable. The core operation of this construction is to use the hardmax to select the value from a specific position. The position encoding in Definition 3.13 satisfied that $p_{i}^{\top}p_{j}<p_{i}^{\top}p_{i}$ for $i\neq j$.
In the first attention head, we use the standard attention, which is not incorporated the incident matrix to clear the column $E$ preparing for writing:
$(W^{(2)}_{K},W^{(2)}_{Q})_{a,b}=\begin{cases}1&\text{if }(a,b)\in\left\{(P_{1},1),(P_{2},2),(B_{\text{global}},2)\right\};\\ 0&\text{otherwise,}\end{cases}\quad(W^{(2)}_{V})_{a,b}=\begin{cases}-1&\text{if }(a,b)=(E,E);\\ 0&\text{otherwise.}\end{cases}$
where $W^{(2)}_{K}$ and $W^{(2)}_{Q}$ are constructed to perform an identity matrix, and $W^{(2)}_{V}$ is constructed to erase the value in column $E$.
We also need another standard attention head to write the value:
$(W^{(1)}_{K},W^{(1)}_{Q})_{a,b}=\begin{cases}1&\text{if }(a,b)\in\{(P_{1},C_{1}),(P_{2},C_{2})\};\\ 0&\text{otherwise,}\end{cases}\quad(W^{(1)}_{V})_{a,b}=\begin{cases}2&\text{if}(a,b)=(D,E)\\ 0&\text{otherwise.}\end{cases},$
where $W^{(1)}_{K}$ and $W^{(1)}_{Q}$ are constructed to indicate the row, and $W^{(1)}_{V}$ is constructed to indicate the column.
Noticing that the writing value head writes values in all rows of column $E$. We establish the MLP layer as follows to erase the unnecessary writing:
$(W^{(1)})_{a,b}$ =$=\begin{cases}1&\text{if }(a,b)=(E,E);\\ -\Omega&\text{if }(a,b)=(B_{\text{global}},E);\\ 0&\text{otherwise,}\end{cases}\quad(W^{(2,3)})_{a,b}=\begin{cases}1&\text{if }(a,b)=(E,E);\\ 0&\text{otherwise,}\end{cases}$
$(W^{(4)})_{a,b}$ =$=\begin{cases}-1&\text{if }(a,b)=(E,E);\\ 0&\text{otherwise,}\end{cases}$
Algorithm 3 Iteration of hyperedge
1:$\mathrm{idx}_{\mathrm{hyperedge}}\leftarrow 0$
[…]
3:$\mathrm{visit}_{\mathrm{hyperedge}},~{}\mathrm{candidates}_{\mathrm{hyperedge}}\leftarrow{\bf 0}_{n_{e}}\|{\bf 1}_{K-1-n_{e}},~{}\Omega\cdot{\bf 1}_{K-1}$
4:$\mathrm{termination}_{\mathrm{hyperedge}}\leftarrow 0$triangleright Initialization of visiting hyperedges
5:
6:procedure VisitHyperedge($\widetilde{A}\in\mathbb{R}_{+}^{n_{v}\times n_{e}}$)
7:     if $\mathrm{termination}_{\mathrm{hyperedge}}=1$ then
8:         $\cdots$ $\triangleright$ Re-initialization of visiting edge, Lemma 4.1
9:         $\mathrm{termination}_{\min}\leftarrow 0$
10:     end if
11:     $\mathrm{idx}_{\mathrm{hyperedge}}\leftarrow\mathrm{idx}_{\mathrm{hyperedge}}+1$triangleright Increment, Lemma 4.2
12:     $\mathrm{val}_{\mathrm{hyperedge}}\leftarrow A[:,\mathrm{idx}_{\mathrm{hyperedge}}]$triangleright Read column from incident matrix, Lemma 4.7
13:     for $i=1\to K-1$ do
14:         $\mathrm{iszero}_{\mathrm{hyperedge}}[i]\leftarrow(\mathrm{val}_{\mathrm{hyperedge}}[i]\leq 0)$triangleright Compare, Lemma 4.4
15:     end for
16:     for $i=1\to K-1$ do
17:         if $\mathrm{iszero}_{\mathrm{hyperedge}}[i]=1$  then
18:              $\mathrm{val}_{\mathrm{hyperedge}}[i]\leftarrow\Omega$triangleright Selection, Lemma 4.1
19:         end if
20:     end for
21:     for $i=1\to K-1$ do
22:         if $\mathrm{val}_{\mathrm{hyperedge}}[i]<\mathrm{candidates}[i]$ then $\triangleright$ Compare, Lemma 4.4
23:              $\mathrm{candidates}[i]\leftarrow\mathrm{val}_{\mathrm{hyperedge}}[i]$triangleright Update variables, Lemma 4.1
24:         end if
25:     end for
26:     $\mathrm{visit}_{\mathrm{hyperedge}}[\mathrm{idx}_{\mathrm{hyperedge}}]\leftarrow 1$triangleright Write scalar to column, Lemma 4.5
27:     $\mathrm{termination}_{\mathrm{hyperedge}}\leftarrow\neg(0~{}\mathrm{in}~{}\mathrm{visit}_{\min})$triangleright Trigger termination, Lemma 4.6
28:     $\mathrm{termination}_{\min}\leftarrow\mathrm{termination}_{\min}\land\mathrm{termination}_{\mathrm{hyperedge}}$triangleright AND, Lemma 4.8
29:end procedure
Algorithm 4 Dijkstra’s algorithm for shortest path
1:procedure HypergraphDijkstra($A\in\mathbb{R}_{+}^{n_{v}\times n_{e}}$, $\mathrm{start}\in\mathbb{N}$)
2:    $\mathrm{prev},~{}\mathrm{dists},~{}\mathrm{dists}_{\mathrm{masked}},~{}\mathrm{changes},~{}\mathrm{iszero}\leftarrow{\bf 0}_{K-1}$
3:    $\mathrm{visit},~{}\mathrm{dists},~{}\mathrm{prev}\leftarrow{\bf 0}_{n_{v}}\|{\bf 1}_{K-1-n_{v}},~{}\Omega\cdot{\bf 1}_{K-1},~{}[K-1]$
4:    $\mathrm{visit}[\mathrm{start}],\mathrm{termination}\leftarrow 0$triangleright Initialization of minimum value and visiting hyperedges
5:
6:    while $\mathrm{termination}$ is $\mathbf{false}$ do
7:         for $i=1\to K-1$ do
8:             if $\mathrm{visit}[i]$ is $\mathbf{true}$ then
9:                 $\mathrm{dists}_{\mathrm{masked}}[i]\leftarrow\Omega$triangleright Selection, Lemma 4.1
10:             else
11:                 $\mathrm{dists}_{\mathrm{masked}}[i]\leftarrow\mathrm{dists}[i]$
12:             end if
13:         end for
14:
15:         GetMinimumValue $(\mathrm{dists}_{\mathrm{masked}})$triangleright Get minimum value, Lemma 5.2
16:
17:         if $\mathrm{termination}_{\min}$ is $\mathbf{true}$ then
18:             $\mathrm{node}\leftarrow\mathrm{idx}_{\mathrm{best}}$
19:             $\operatorname{dist}\leftarrow\mathrm{val}_{\mathrm{best}}$triangleright Selection, Lemma 4.1
20:         end if
21:
22:         VisitHyperdege $(A)$triangleright Degradation, Theorem 6.1
23:
24:         if $\mathrm{termination}_{\mathrm{hyperedge}}=1$ then
25:             $\mathrm{candidates}\leftarrow\mathrm{candidates}_{\mathrm{hyperedge}}$triangleright Selection, Lemma 4.1
26:         end if
27:         for $i=1\to K-1$ do
28:             $\mathrm{candidates}[i]\leftarrow\mathrm{candidates}[i]+\operatorname{dist}$triangleright Addition, Lemma 4.10
29:         end for
30:         for $i=1\to K-1$ do
31:             $\mathrm{changes}[i]\leftarrow\mathrm{candidates}[i]<\mathrm{dists}[i]$triangleright Compare, Lemma 4.4
32:         end for
33:         for $i=1\to K-1$ do
34:             if $\mathrm{termination}_{\min}=0$ and $\mathrm{iszero}[i]=1$ then
35:                 $\mathrm{changes}[i]\leftarrow 0$triangleright repeat AND, Lemma 4.9
36:             end if
37:         end for
38:         for $i=1\to K-1$ do
39:             if $\mathrm{changes}[i]=1$ then
40:                 $\mathrm{prev}[i],~{}\mathrm{dists}[i]\leftarrow\mathrm{node},\mathrm{candidates}[i]$triangleright Selection. Lemma 4.1
41:             end if
42:         end for
43:         $\mathrm{visit}[\mathrm{node}]\leftarrow\mathrm{visit}[\mathrm{node}]+\mathrm{termination}_{\min}$triangleright Write scalar to column, Lemma 4.5
44:         $\mathrm{termination}\leftarrow\neg(0~{}\mathrm{in}~{}\mathrm{visit})$triangleright Trigger termination, Lemma 4.6
45:    end while
46:    return $\mathrm{prev},~{}\mathrm{dists}$
47:end procedure
Algorithm 5 Helly, Algorithm 3 in [5]
1:$\mathrm{idx}_{x}\leftarrow 0$
2:$\mathrm{idx}_{y}\leftarrow 1$
3:$\mathrm{idx}_{v}\leftarrow 0$
4:$\mathrm{termination}\leftarrow 0$
5:$\mathrm{helly}\leftarrow 1$triangleright Initialization of Helly
6:
7:while $\mathrm{termination}=0$ do
8:     $\mathrm{hyperedge}_{x}\leftarrow A[:,\mathrm{idx}_{x}]$
9:     $\mathrm{hyperedge}_{y}\leftarrow A[:,\mathrm{idx}_{y}]$
10:     $\mathrm{hyperedge}_{v}\leftarrow A[:,\mathrm{idx}_{v}]$triangleright Read column from incident matrix, Lemma 4.7
11:     $\mathrm{hyperedge}_{x}\leftarrow\mathrm{hyperedge}_{x}>0$
12:     $\mathrm{hyperedge}_{y}\leftarrow\mathrm{hyperedge}_{y}>0$
13:     $\mathrm{hyperedge}_{v}\leftarrow\mathrm{hyperedge}_{v}>0$triangleright Compare, Lemma 4.4
14:     $\mathrm{intersection}\leftarrow\mathrm{hyperedge}_{x}\wedge\mathrm{hyperedge}_{y}\wedge\mathrm{hyperedge}_{v}$triangleright AND, Lemma 4.8
15:     $\mathrm{helly}_{v}\leftarrow\neg$(1 in $\mathrm{intersection}$)
16:     if $\mathrm{helly}_{v}=1$ then
17:         $\mathrm{helly}\leftarrow 0$
18:         $\mathrm{termination}\leftarrow 1$triangleright Selection, Lemma 4.1
19:     end if
20:     $\mathrm{idx}_{v}\leftarrow\mathrm{idx}_{v}+1$triangleright Increment, Lemma 4.2
21:     if $\mathrm{idx}_{v}>n_{v}$ then
22:         $\mathrm{idx}_{v}\leftarrow 1$triangleright Selection, Lemma 4.1
23:         $\mathrm{idx}_{y}\leftarrow\mathrm{idx}_{y}+1$triangleright Selection and Increment, Lemma 4.1 and Lemma 4.2
24:     end if
25:     if $\mathrm{idx}_{y}>n_{v}$ then
26:         $\mathrm{idx}_{x}\leftarrow\mathrm{idx}_{x}+1$
27:         $\mathrm{idx}_{y}\leftarrow\mathrm{idx}_{x}+1$triangleright Selection and Increment, Lemma 4.1 and Lemma 4.2
28:     end if
29:     if $\mathrm{idx}_{x}=n_{v}$ then
30:         $\mathrm{termination}\leftarrow 1$triangleright Selection, Lemma 4.1
31:     end if
32:end while
33:return $\mathrm{helly}$
4.4 Comparison
Lemma 4.4 (Comparison).
If the following conditions hold:
  • Let the transformer be defined as Definition 3.11.
  • Let $\Omega$ represent the maximum absolute value within a clause.
  • Let $C,~{}D$ be the index for the column for comparing.
  • Let $E$ be the index for the target column to write the comparison result.
Then, we can show that a single-layer transformer can simulate a comparison operation, which achieves writing $X[:,E]\leftarrow X[:,C]<X[:,D]$.
Proof.
Because only the MLP layer is needed, we set all the parameters in the attention layer to 0 while keeping the residual connection. Let $S_{1},~{}S_{2}$ be the index for the scratchpad. We construct the weights in MLP as follows:
$(W^{(1)})_{a,b}$ =$=\begin{cases}1&\text{if }(a,b)\in\{(E,E),(D,S_{1}),(D,S_{2})\};\\ -1&\text{if }(a,b)\in\{(C,S_{1}),(C,S_{2})\};\\ -\Omega^{-1}&\text{if }(a,b)\in\{(B_{\mathrm{global}},S_{2}),(B_{\mathrm{local}},S_{2})\};\\ 0&\text{otherwise,}\end{cases}\quad$ $(W^{(2)})_{a,b}=\begin{cases}1&\text{if }(a,b)=(E,E);\\ \Omega&\text{if }(a,b)=(S_{1},S_{1});\\ -\Omega&\text{if }(a,b)=(S_{2},S_{1});\\ 0&\text{otherwise.}\end{cases}$
$(W^{(3)})_{a,b}$ =$=\begin{cases}1&\text{if }(a,b)\in\{(E,E),(S_{1},S_{1})\};\\ 0&\text{otherwise,}\end{cases}\quad$ $(W^{(4)})_{a,b}=\begin{cases}-1&\text{if }(a,b)=(E,E);\\ 1&\text{if }(a,b)=(S_{1},E);\\ 0&\text{otherwise.}\end{cases},$
where $W^{(1)}$ and $W^{(2)}$ are defined to simulate less than function, $W^{(3)}$ is defined as identity layer, $W^{(4)}$ is defined to erase the previous value in $X[:,E]$ and write the term on scarctchpad to column $E$. ∎
4.5 Write Scalar to Column
Lemma 4.5 (Write scalar to column).
If the following conditions hold:
  • Let the transformer be defined as Definition 3.11.
  • Let $P$ be the index for the position embeddings.
  • Let $\Omega$ represent the maximum absolute value within a clause.
  • Let $D$ be the index for source value, i.e., source value is $X[0,D]$.
  • Let $C_{1},C_{2}$ be the position embedding for target row, $E$ be the index for target column.
Then, we can show that a single-layer transformer can simulate the operation of writing a value to a column, i.e., $X[C,E]\leftarrow X[0,D]$.
Proof.
The key purpose is to move the information from the global variable to the local variable. The core operation of this construction is to use hardmax to select the value from a specific position. The position encoding in Definition 3.13 satisfied that $p_{i}^{\top}p_{j}<p_{i}^{\top}p_{i}$ for $i\neq j$. Since the MLP layer is not needed, we set all parameters to 00.
First, we construct the first attention layer to write scalar:
$(W^{(1)}_{K},W^{(1)}_{Q})_{a,b}=\begin{cases}1&\text{if }(a,b)\in\{(P_{1},C_{1}),(P_{2},C_{2})\};\\ 0&\text{otherwise,}\end{cases}\quad(W^{(1)}_{V})_{a,b}=\begin{cases}2&\text{if}(a,b)=(D,E);\\ 0&\text{otherwise.}\end{cases}$
where $(W^{(1)}_{K}$ and $W^{(1)}_{Q})$ are constructed to find the row $C$, and $W^{(1)}_{V})$ is constructed to write scalar from column $D$ to column $E$.
The above construction will write some unwanted value to the top row, so we construct another attention head to erase the unwanted value:
$(W^{(2)}_{K},W^{(2)}_{Q})_{a,b}$ =$=\begin{cases}1&\text{if }(a,b)\in\left\{(P_{1},1),(P_{2},2),(B_{\text{global}},2)\right\}\\ 0&\text{otherwise,}\end{cases}$
$(W^{(2)}_{V})_{a,b}$ =$=\begin{cases}-1&\text{if }(a,b)=(B_{\text{global}},E);\\ 0&\text{otherwise.}\end{cases}$
where we use $B_{\mathrm{global}}$ is used to store global bias. ∎
4.6 Termination
Lemma 4.6 (Termination).
If the following conditions hold:
  • Let the transformer be defined as Definition 3.11.
  • Let $P$ be the index for the position embeddings.
  • Let $\Omega$ represent the maximum absolute value within a clause.
  • Let $C$ be the index for the executed variable.
  • Let $E$ be the index for the target column.
Then we can show that a single-layer transformer can simulate the operation of writing a value to a column, i.e. $X[1,E]\leftarrow(\mathrm{no~{}zero~{}in~{}}X[2:,C])$.
Proof.
We construct the first attention layer to erase the previous value in column $E$:
$(W^{(1)}_{K},W^{(1)}_{Q})_{a,b}=\begin{cases}1&\text{if }(a,b)\in\left\{(P_{1},1),(P_{2},2),(B_{\text{global}},2)\right\};\\ 0&\text{otherwise,}\end{cases}\quad(W^{(1)}_{V})_{a,b}=\begin{cases}-1&\text{if }(a,b)=(E,E);\\ 1&\text{if }(a,b)=(B_{\text{global}},E);\\ 0&\text{otherwise,}\end{cases}$
where $W^{(1)}_{K}$ and $W^{(1)}_{Q}$ are constructed as identity matrix, $W^{(1)}_{V}$ is constructed to replace the previous value in column $E$ by $1$, which will be used in the following construction.
In the second attention head, we want to construct an attention matrix that can extract information from all the entries of $X[2:,C]$:
$(W^{(2)}_{K})_{a,b}$ =$=\begin{cases}-1&\text{if }(a,b)\in\{(B_{\text{global}},1),(B_{\text{global}},2)\};\\ 1&\text{if }(a,b)\in\{(B_{\text{local}},1),(B_{\text{local}},2)\};\\ 0&\text{otherwise,}\end{cases}$
$(W^{(2)}_{Q})_{a,b}$ =$=\begin{cases}1&\text{if }(a,b)\in\{(B_{\text{global}},1),(B_{\text{global}},2)\};\\ -1&\text{if }(a,b)\in\{(B_{\text{local}},1),(B_{\text{local}},2)\};\\ 0&\text{otherwise,}\end{cases}$
and the attention matrix is as presented below:
$\sigma\!\left(XW^{(2)}_{Q}W^{(2)\top}_{K}X^{\top}\right)={\sigma\left(2\left[\begin{array}[]{c|cccccc}-1&1&\cdots&1\\ \hline\cr 1&-1&\cdots&-1\\ \vdots&\vdots&\ddots&\vdots\\ 1&-1&\cdots&-1\end{array}\right]\right)}={\left[\begin{array}[]{c|ccccc}0&\frac{1}{n}&\cdots&\frac{1}{n}\\ \hline\cr 1&0&\cdots&0\\ \vdots&\vdots&\ddots&\vdots\\ 1&0&\cdots&0\end{array}\right]}.$
We construct the value matrix as:
$(W^{(2)}_{V})_{a,b}=\begin{cases}\Omega&\text{if }(a,b)=(C,E);\\ -\Omega&\text{if }(a,b)=(B_{\text{local}},E);\\ 0&\text{otherwise,}\end{cases}$
after this, the top entry of column $E$ is $1-\Omega+\frac{\Omega}{n}\sum_{i}(X[i,C])$. Applying ReLU units, we construct the MLP layer as follows:
$(W^{(1)})_{a,b}$ =$=\begin{cases}1&\text{if }(a,b)\in\{(E,E),(E,S_{1})\}\\ -1&\text{if }(a,b)=(E,S_{2});\\ 0&\text{otherwise,}\end{cases}\quad(W^{(2,3)})_{a,b}=\begin{cases}1&\text{if }(a,b)\in\{(E,E),(S_{1},S_{1}),(S_{2},S_{2})\};\\ 0&\text{otherwise,}\end{cases}$
$(W^{(4)})_{a,b}$ =$=\begin{cases}1&\text{if }(a,b)\in\{(E,E),(S_{2},E)\};\\ -1&\text{if }(a,b)=(S_{1},E);\\ 0&\text{otherwise.}\end{cases}$
which writes $\phi(1-\Omega+\frac{\Omega}{n}\sum_{i}(X[i,C]))$ to $X[1,E]$. It’s easy to know that if only if all the value are $1$ in $X[2:,C]$, $X[1,E]$ gets value $1$. ∎
4.7 Read from Incident Matrix
Lemma 4.7 (Read from incident matrix).
If the following conditions hold:
  • Let the transformer be defined as Definition 3.11.
  • Let $P$ be the index for the position embeddings.
  • Let $C$ be the index for the source row/column index.
  • Let $D$ be the index for the target column.
  • Let $P_{\mathrm{cur}}$ be the index for the position embedding of row/column $C$.
Then we can show that:
  • Part 1. A single-layer transformer can simulate the operation of reading a row from $A$, i.e. $X[:,D]\leftarrow\widetilde{A}[C,:]$.
  • Part 2. A single-layer transformer can simulate the operation of reading a column from $A$, i.e. $X[:,D]\leftarrow\widetilde{A}[:,C]$.
Proof.
Following from Definition 3.6, we can either employ $\psi^{(i)}(X,\widetilde{A})$ or $\psi^{(i)}(X,\widetilde{A}^{\top})$. So, reading a column and reading a row is equivalent in our setting. Considering the reading row case, we construct one attention head as follows:
$(W_{K}^{(1)},W_{Q}^{(1)})_{a,b}$ =$=\begin{cases}1&\text{if }(a,b)\in\{(P_{1},(P_{\text{cur}})_{1}),(P_{2},(P_{\text{cur}})_{2})\};\\ 0&\text{otherwise,}\end{cases}$
$(W_{V}^{(1)})_{a,b}$ =$=\begin{cases}2&\text{if }(a,b)=(B_{\text{global}},D);\\ 0&\text{otherwise,}\end{cases}$
where $W_{Q}^{(1)}$ and $W_{K}^{(1)}$ are constricuted to get row $C$ following from Definition 3.13. $W_{V}^{(1)}$ is used to move row $C$ of matrix $A$ to column $D$ of matrix $X$.
For the second attention head, we use a standard attention head to erase the previous value in column $D$.
$(W^{(2)}_{K},W^{(2)}_{Q})_{a,b}=\begin{cases}1&\text{if }(a,b)\in\{(P_{1},1),(P_{2},2),(B_{\text{global}},2)\};\\ 0&\text{otherwise,}\end{cases}\;(W^{(2)}_{V})_{a,b}=\begin{cases}-1&\text{if }(a,b)=(D,D);\\ 0&\text{otherwise,}\end{cases}$
Where $W^{(2)}_{K},~{}W^{(2)}_{Q}$ are constructed to make the attention matrix as identity matrix, $W^{(2)}_{V}$ is constructed to erase value. For the MLP layer, we just make it as the residential connection by setting all parameters to 00. ∎
4.8 AND
Lemma 4.8 (AND).
If the following conditions hold:
  • Let the transformer be defined as Definition 3.11.
  • Let $\Omega$ represent the maximum absolute value within a clause.
  • Let $C$ and $D$ be the index for the executed variable.
  • Let $E$ be the index for the target column.
Then we can show that a single-layer transformer can simulate the operation of AND, i.e. $X[1,E]\leftarrow X[1,C]\wedge X[1,D]$.
Proof.
In this construction, we want to have $\phi(X[1,C]+X[1,D]-1)$, so that if and only if $X[1,C]=1$ and $X[1,D]=1$, we have $\phi(X[1,C]+X[1,D]-1)=1$, otherwise $\phi(X[1,C]+X[1,D]-1)=0$. Because only the MLP layer is needed, we set all the parameters in the attention layer to 0 while keeping the residual connection. We construct MLP layers as follows:
$(W^{(1)})_{a,b}$ =$=\begin{cases}1&\text{if }(a,b)\in\{(E,E),(C,S),(D,S)\};\\ -1&\text{if }(a,b)=(B_{\mathrm{global}},S);\\ 0&\text{otherwise,}\end{cases}\quad(W^{(2,3)})_{a,b}=\begin{cases}1&\text{if }(a,b)\in\{(E,E),(S,S)\};\\ 0&\text{otherwise,}\end{cases}$
$(W^{(4)})_{a,b}$ =$=\begin{cases}-1&\text{if }(a,b)=(E,E);\\ 1&\text{if }(a,b)=(S,E);\\ 0&\text{otherwise,}\end{cases}$
where $W^{(1)}$ is the core of the construction, $W^{(2,3)}$ are defined as an identity operator, $W^{(4)}$ is used to erase the previous value and move the result in scratchpad to column $E$.
4.9 Repeat AND
Lemma 4.9 (Repeat AND).
If the following conditions hold:
  • Let the transformer be defined as Definition 3.11.
  • Let $\Omega$ represent the maximum absolute value within a clause.
  • Let $C$ be the index for a Boolean value.
  • Let $D$ and $E$ be the index for two columns, where each entry is a Boolean value.
Then we can show that a single-layer transformer can simulate the operation of repeat AND, i.e., $X[i,D]\leftarrow X[1,C]\wedge X[i,D]\wedge\neg X[i,E]$ for $i\in\{2,3,\cdots,K\}$.
Proof.
In this construction, we only use MLP layers. For multi-head attention layers, we just make it as the residential connection by setting all parameters to 00.
$(W^{(1)})_{a,b}$ =$=\begin{cases}1&\text{if }(a,b)\in\{(D,D),(C,D),(D,S_{1})\};\\ -1&\text{if }(a,b)\in\{(E,D),(B_{\text{global}},D)(B_{\text{local}},D)\};\\ 0&\text{otherwise,}\end{cases}$
$(W^{(2,3)})_{a,b}$ =$=\begin{cases}1&\text{if }(a,b)\in\{(D,D),(S_{1},S_{1})\};\\ 0&\text{otherwise,}\end{cases}\quad(W^{(4)})_{a,b}=\begin{cases}1&\text{if }(a,b)=(D,D);\\ -1&\text{if }(a,b)=(S_{1},D);\\ 0&\text{otherwise.}\end{cases}$
where $(W^{(1)})$ is used to construct $\phi(X[1,C]+X[i,D]-X[i,E]-1)$, $(W^{(2,3)})$ are constructed as identity layers, $(W^{(4)})$ is constructed to erase previous value in column $D$. ∎
4.10 Repeat Addition
Lemma 4.10 (Repeat addition).
If the following conditions hold:
  • Let the transformer be defined as Definition 3.11.
  • Let $\Omega$ represent the maximum absolute value within a clause.
  • Let $P$ be the index for the position embeddings.
  • Let $C$ be the index for a scalar.
  • Let $D$ be the index for a column.
Then we can show that a single-layer transformer can simulate the operation of repeat addition, i.e. $X[:,D]\leftarrow{\bf 1}_{K}\cdot X[1,C]+X[:,D]$.
Proof.
We build the first attention head as:
$(W^{(1)}_{K})_{a,b}$ =$=\begin{cases}1&\text{if }(a,b)\in\{(B_{\text{global}},1),(B_{\text{global}},2)\};\\ 0&\text{otherwise,}\end{cases}$
$(W^{(1)}_{Q})_{a,b}$ =$=\begin{cases}1&\text{if }(a,b)\in\left\{(B_{\text{global}},1),(B_{\text{global}},2),(B_{\text{local}},1),(B_{\text{local}},2)\right\};\\ 0&\text{otherwise,}\end{cases}$
$(W^{(1)}_{V})_{a,b}$ =$=\begin{cases}1&\text{if }(a,b)=(C,D);\\ 0&\text{otherwise,}\end{cases}$
where we have
$\sigma\!\left(XW^{(1)}_{Q}W^{(1)\top}_{K}X^{\top}\right)={\sigma\left(2\left[\begin{array}[]{c|cccccc}1&0&\cdots&0\\ \hline\cr 1&0&\cdots&0\\ \vdots&\vdots&\ddots&\vdots\\ 1&0&\cdots&0\end{array}\right]\right)}={\left[\begin{array}[]{c|ccccc}1&0&\cdots&0\\ \hline\cr 1&0&\cdots&0\\ \vdots&\vdots&\ddots&\vdots\\ 1&0&\cdots&0\end{array}\right]}.$
Using the first attention head, we can repeat $X[1,C]$ as a column. For the second attention head, we just select the column to erase the first value in column $D$.
$(W^{(2)}_{K},W^{(2)}_{Q})_{a,b}=\begin{cases}1&\text{if }(a,b)\in\{(P_{1},1),(P_{2},2),(B_{\text{global}},2)\};\\ 0&\text{otherwise,}\end{cases}\;(W^{(2)}_{V})_{a,b}=\begin{cases}-1&\text{if }(a,b)=(C,D);\\ 0&\text{otherwise.}\end{cases}$
Where $W^{(2)}_{K},~{}W^{(2)}_{Q}$ are constructed to make the attention matrix as an identity matrix, $W^{(2)}_{V}$ is constructed to erase value. For the MLP layer, we just make it as the residential connection by setting all parameters to 00. ∎
5 Simulation
We present the simulation result of visiting hyperedges iteratively in Section 5.1. We discuss the result of getting a minimum value in Section 5.2. We discuss the simulation result of Dijkstra’s Algorithm in Section 5.3.
5.1 Iteration of Visiting Hyperedges
Lemma 5.1 (Visiting hyperedges iteratively).
A looped transformer $h_{T}$ exists, where each layer is defined as in Definition 3.11, consisting of 10 layers where each layer includes 3 attention heads with feature dimension of $O(1)$. This transformer simulates the operation of iteratively visiting hyperedges for weighted hypergraphs, accommodating up to $O(\widehat{\delta}^{-1})$ vertices and $O(\widehat{\delta}^{-1})$ hyperedges.
Proof.
Let the operation of visiting hyperedges iteratively be defined as described in Algorithm 3. This operation requires 1 increment operation, which requires single layer transformer to construct the following from Lemma 4.2, 2 compare operations, which require 3 layers transformer to construct the following from Lemma 4.4, 3 selection operations, which require 3 layers transformer to construct following from Lemma 4.1, 1 AND operation which require single layer transformer to construct following from Lemma 4.8, 1 read-from-incident-matrix operation which requires single layer transformer to construct following from Lemma 4.7, 1 write-scalar-to-column operation which requires single layer transformer to construct following from Lemma 4.5, and 1 trigger-termination operation which require single layer transformer to construct following from Lemma 4.6. The whole algorithm can be constructed by
$1+1+1+1+1+2+3=10$
layers of the transformer. ∎
5.2 Iteration of Minimum Value
Lemma 5.2 (Get minimum value).
If the following conditions hold:
  • Let the transformer be defined as Definition 3.11.
Then, we can show that a 7-layer transformer can simulate the operation of getting the minimum value in Algorithm 2.
Proof.
Following from Lemma 4.1, Lemma 4.2, Lemma 4.3, Lemma 4.4, Lemma 4.5, and Lemma 4.6, we can perform all the operation of Algorithm 2, thus we finish the proof. ∎
5.3 Dijkstra’s Algorithm
Theorem 5.3 (Dijkstra’s Algorithm on hypergraph).
A looped transformer $h_{T}$ exists, where each layer is defined as in Definition 3.11, consisting of 27 layers, where each layer includes 3 attention heads with feature dimension of $O(1)$. This transformer simulates Dijkstra’s Algorithm iteratively for weighted hypergraphs, supporting up to $O(\widehat{\delta}^{-1})$ vertices and $O(\widehat{\delta}^{-1})$ hyperedges.
Proof.
Let Dijkstra’s algorithm be considered as described in Algorithm 4. Following from Lemma 5.1 and Lemma 5.2, the operation of iteratively visiting vertices and hyperedges requires 18 layers, as established in these lemmas. For the remaining part of the algorithm, the operations include 4 selection operations, 1 add operation, 1 compare operation, 1 repeat AND operation, 1 write scalar to column operation, and 1 trigger termination operation. According to Lemma 4.1, Lemma 4.10, Lemma 4.9, Lemma 4.4, Lemma 4.5, and Lemma 4.6, these operations together require 9 layers. Therefore, the total number of layers required to simulate Dijkstra’s algorithm is:
$18+9=27$
layers of the transformer. ∎
6 Main Results
We are now ready to show our main results based on our previous components.
6.1 Degradation
We observe that [13] presents the running results of several algorithms on graphs, which primarily include Dijkstra’s algorithm for the shortest path, Breadth-First Search (BFS), and Depth-First Search (DFS). [13] provides a general representation of such algorithms in Algorithm 7 of their paper. Given that Dijkstra’s algorithm for the shortest path, BFS, and DFS can be straightforwardly extended to hypergraphs [21] by identifying the shortest hyperedge between two nodes and treating it as the distance between those nodes, we propose a degradation mechanism in Algorithm 3. This mechanism allows dynamic access to the shortest hyperedge between two vertices while maintaining the feature dimension of the Looped Transformer at $O(1)$. Using this mechanism, we can extend Dijkstra’s algorithm for the shortest path, BFS, and DFS from graphs to hypergraphs. We formally state the theorem regarding the degradation mechanism as follows:
Theorem 6.1 (Degradation).
A looped transformer $h_{T}$ exists, where each layer is defined as in Definition 3.11, consisting of 10 layers, where each layer includes 3 attention heads with feature dimension of $O(1)$. This transformer simulates the degradation operation, i.e., the degradation of the incident matrix to the adjacency matrix while retaining only the shortest hyperedge between two vertices, for weighted hypergraphs, supporting up to $O(\widehat{\delta}^{-1})$ vertices and $O(\widehat{\delta}^{-1})$ hyperedges.
Proof.
Using the same construction as Lemma 5.1, we finish the proof. ∎
6.2 Helly
The Helly property in a hypergraph refers to a specific condition in the family of its hyperedges. A hypergraph is said to satisfy the Helly property if, for every collection of its hyperedges, whenever the intersection of every pair of hyperedges in the collection is non-empty, there exists at least one hyperedge in the collection that intersects all the others. In other words, if the pairwise intersections of a set of hyperedges are non-empty, there must be a single hyperedge that intersects every hyperedge in the set. This property is named after Helly’s theorem in convex geometry, which inspired its application in combinatorial settings like hypergraphs.
We have reformulated the algorithm for determining whether a hypergraph possesses the Helly property, as presented in Algorithm 3 of [5], into a form that the Looped Transformer can execute. This reformulation is represented by our Algorithm 5. We now state the following theorem:
Theorem 6.2 (Helly).
A looped transformer $h_{T}$ exists, where each layer is defined as in Definition 3.11, consisting of 11 layers, where each layer includes 3 attention heads with feature dimension of $O(1)$. This transformer simulates the Helly algorithm (Algorithm 5) for weighted hypergraphs, handling up to $O(\widehat{\delta}^{-1})$ vertices and $O(\widehat{\delta}^{-1})$ hyperedges.
Proof.
Let the degradation operation be defined as described in Algorithm 3. This operation requires 3 increment operations, 5 selection operations, 1 AND operation, 1 compare operation, and 1 read-from-incident-matrix operation. By applying Lemma 4.2, Lemma 4.1, Lemma 4.8, Lemma 4.4, and Lemma 4.7, each of these operations can be constructed using a single-layer transformer, as detailed in the corresponding lemmas. The total number of layers required for the entire degradation operation is:
$3+5+1+1+1=11$
Thus, the degradation operation can be constructed using 11 layers of transformer. This completes the proof. ∎
7 Conclusion
In this work, we extended the capabilities of Loop Transformers to the domain of hypergraphs, addressing the computational challenges posed by their complex structures. By introducing a degradation mechanism to simulate graph-based algorithms on hypergraphs and a hyperedge-aware encoding scheme for hypergraph-specific algorithms, we demonstrated the feasibility of using Transformers for hypergraph algorithm simulation. Our results, supported by theoretical guarantees, underscore the potential of Loop Transformers as general-purpose computational tools capable of bridging neural networks and combinatorial optimization tasks on structured, high-dimensional data. These findings not only expand the applicability of Transformers but also open new avenues for solving real-world problems.